OpenStudy (s):

(x^2)*y'' + 5*x*y' - 5y = 0

OpenStudy (s):

Im not sure what happens to the x.. do u ignore it when u do r^2 + 5r - 5 = 0 or u do something else?

OpenStudy (anonymous):

this is called Cauchy-Euler equation. because of it shapes , it is very easy to solve it y=t^r y'= r t^(r-1) y''= (r(r-1))t^(r-2) r(r-1) + 5 r -5 =0 r^2 - r+ 5r-5=0 r^2+4r-5=0 (r +5)(r -1) r=-5,1

OpenStudy (s):

Ooooo ok, i remember now thanks!